# MT2ST: Adaptive Multi-Task to Single-Task Learning

### v2

@@ Abstract
The conventional training approaches often face challenges in balancing the breadth of multi-task learning (MTL) with the depth of single-task learning (STL). To address this issue, we introduce the Multi-Task to Single-Task (MT2ST) framework, a groundbreaking approach that can combine the generalizability of MTL with the precision of STL. Our work include two strategies: ’Diminish’ and ’Switch’. ’Diminish’ Strategy will gradually reduce the influence of auxiliary tasks, while the ’Switch’ strategy involves a shift from multi-tasking to single-tasking at a specific timepoint at the training process.
In this paper, we propose the Multi-Task to Single-Task (MT2ST) framework, a novel approach that significantly enhances the efficiency and accuracy of word embedding training while concurrently addressing prevalent issues such as overfitting. Our empirical studies demonstrate that MT2ST can reduce training time by 67% when contrasted with single-task learning approaches, and by 13% compared to traditional multi-task learning methods. These findings underscore MT2ST’s potential to be a powerful tools for word embedding training acceleration. The code implementation is can be found at: https://github.com/NoakLiu/MT2ST-Word-Embeddings-Acceleration.
@@
Efficient ML, ML Acceleration
@@ I Introduction
Recent developments in natural language processing have highlighted that word embeddings are foundational to many NLP tasks,its performance and training efficiency directly impacting the effectiveness of linguistic tasks.
Traditional word embedding training methods predominantly rely on either MTL, which offers a wide learning scope at the expense of task-specific detail, or STL, which excels in focus on specific task but lacks the comprehensiveness of MTL. [ZWX+24]. Our research introduces a novel paradigm, the MT2ST framework, which combines comprehensive scope of MTL and task-sepecific precision of STL together. [CPKC09]
The MT2ST framework incorporates two specific strategies for training word embeddings: ’Diminish’ and ’Switch’. The Diminish strategy innovatively reduces the impact of auxiliary tasks over time, ensuring a gradual shift to STL. In contrast, the ’Switch’ strategy splits training into two stages: starting with MTL and then making a direct transition to STL. MT2ST framework has been proven effectiveness in our emprical study. [BOHG19]
Our contributions can be summarized as follows:
  • Bridge the gap between MTL and STL with the unique MT2ST framework.
  • Introduce two pioneering strategies within this framework. ’Diminish’ and ’Switch’.
  • Empirically demonstrate MT2ST’s superiority in training efficiency and model performance. [SK18]
@@ II Motivation
II-A Challenges in Single-Task Word Embedding Training
Word embeddings map words to continuous vector spaces, reflecting their semantic meanings. It has wide applications in text classification, sentiment analysis, and machine translation. [AX23] However, its training process is hindered by several constraints, which are as follows:
@@ II-A Challenges in Single-Task Word Embedding Training
Data Volume Requirement: The need for large-scale textual datasets to train word embeddings results in significant time and computational resource consumption. Processing these large datasets is inherently slow and demands extensive computational power. [EWVD19]
Computational Complexity: Training involves complex mathematical computations. These algorithms, while effective, require high computational resources, leading to increased training duration and processing power needs. Some recent work such as [LWJV24], [LJP24] has proven the importance of efficient computing in reducing computational resources through algorithm and system co-design optimization.
Optimization Difficulties: Optimizing word embedding models for accurate semantic representation is a meticulous process. It often necessitates iterative tuning and adjustments, which are resource-intensive and time-consuming.
II-B Optimizing Training Efficiency in NLP Models with Multi-Task Learning
The training efficiency of word embeddings is hindered by the mismatch between large training datasets and narrow objective function metrics. This often leads to underutilization of datasets and focuses on limited semantic or syntactic aspects, causing inefficiencies. Altough work like attention based multi-task learning [LJD19] can adjust weights for different task overtim, models still struggle with polysemous words due to objective functions missing their subtle meanings, leading to repetitive training and reduced efficiency.
Multi-task learning (MTL) emerges as an effective solution for these inefficiencies. [ZYY+23] Recent research indicates that training on multiple related tasks can yield computational benefits. MTL has wide application in industry, such as building end-to-end dialogue modelling. [SSM+22] MTL enables models to share core representations across tasks, enhancing word embeddings with a richer semantic understanding, thus improving efficiency.
II-C Potential Inefficiency of Multi-Task Learning in Word Embedding Training
MTL brings notable benefits but can also introduce inefficiencies, especially in word embedding training. In MTL, the shared embedding layer across tasks may receive divergent gradient directions from different tasks, which may conflict.
This issue arises as tasks emphasize distinct aspects of word meanings. For example, sentiment analysis prioritizes emotional connotations, while part-of-speech tagging focuses on grammatical features. Consequently, the embedding layer faces mixed directions of gradient descent in tuning embeddings, which can potentially result in less effective task-specific embeddings. [LLJ+21]
Additionally, computational demands rise in MTL. Each task’s will update the parameters the embedding layer, which leads to time inefficiency and high resource usage. This issue intensifies with large vocabularies and high-dimensional embeddings common in NLP.
@@ III-A MT2ST Framework
In this research, we propose ”MT2ST” (Multi-Task to Single-Task) model to improve the efficiency of word embedding training, specifically addressing the challenges in multi-task learning (MTL) environments.
The process begins with a multi-task learning phase, where a unified model, featuring a shared embedding layer, is simultaneously trained across various NLP tasks. This phase utilizes the advantages of MTL, enabling the model to absorb a wide array of linguistic features and nuances. The shared embedding layer benefits from varied linguistic inputs, fostering a more comprehensive word representation.
Once the multi-task phase is established, MT2ST shifts focus to single-task learning. Here, the pre-trained shared embeddings are fine-tuned in a task-specific manner. This targeted fine-tuning aligns the embeddings more closely with the distinct demands and subtleties of each task, ensuring that while they maintain the general linguistic knowledge from the MTL phase, they are also tailored for superior performance in specific tasks.
MT2ST employs advanced techniques for a smooth and effective transition from multi-task to single-task learning, which provides a fine-tuning and pretraining method for effcient NLP[TLJ+23]. These include adaptive learning rate mechanisms and selective freezing of certain embedding vector dimensions based on their relevance to the specific task. These strategies enable a balance between retaining the broad linguistic understanding gained in the MTL phase and focusing on task-specific details during the single-task phase.
@@ III-B Model Construction
Instead of mono-task fine-tuning used in Multi-Task Supervised Pretraining [MST+20], our MT2ST model develop an adaptive method to accelerate word embedding training within a multi-task to single-task learning framework.
@@ III-B2 Embedding Layer
  • Learns dense vector representations of words, capturing semantic and syntactic information.
@@ III-C Model Overview
Figure 1: MT2ST Word Embedding Training Framework Overview22footnotemark: 2
@@ III-D MT2ST: Diminish Strategy
Figure 2: MT2ST-Dimish: Dimish Strategy for Auxiliary Tasks in Training to Single Task Learning
@@
Figure 3: MT2ST to realize quicker convergence [An example of MT2ST Diminish]
@@ III-E MT2ST: Switch Strategy
Figure 4: MT2ST-Switch: Optimal Solution Area Searching and Task Switching
Post $T_{switch}$, the update focuses solely on the primary task:
@@ IV Result Analysis
To evaluate the performance of MT2ST, various experimental setups were implemented (experimental settings are available in the appendix).
In the first experimental group, the approach was to directly fine-tune the pre-pretrained embeddings on a dataset dedicated to a single task. Observations from this experiment indicated a steady decrease in Loss, but there were fluctuations in accuracy, indicating overfitting in the model.
Figure 5: Accuracy and Loss change of Original Mode
In the second experimental series, a multi-task learning approach was applied. This approach led to a consistent improvement in both Loss and accuracy metrics. The model achieved a notable accuracy rate of 68% on the test set.
Figure 6: Loss change (Multi-Task Learning)
Figure 7: Accuracy change (Multi-Task Learning)
In the third set of experiments, the MT2ST-diminish strategy was employed. Unlike previous tests, the Loss did not decrease uniformly, which was attributed to the reduction of auxiliary task Loss during the training process. However, the model’s accuracy consistently improved, ultimately achieving a higher test set accuracy of 71% compared to the standard multi-task learning approach. This result highlights the effectiveness of the MT2ST-diminish strategy in facilitating a smooth transition from multi-task to task-specific learning, thereby enhancing the model’s overall performance.
Figure 8: Loss change in Mode (MT2ST: Dimish)
Figure 9: Accuracy change (MT2ST: Dimish)
The fourth experimental group implemented fine-tuning on the dataset using predefined embeddings through the MT2ST-switch strategy. This approach consisted of two phases: initial multi-task learning followed by task-specific learning, with an abrupt transition between the two. As indicated by the red line in the graph, a step change in Loss occurred after the switch, yet there was a steady improvement in both phases. This method achieved a high score of 77% on the test set, demonstrating its effectiveness.”
Figure 10: Loss change (MT2ST: Switch)
Figure 11: Accuracy change (MT2ST: Swtich)
@@ V Discussion
Our empirical experiments reveal that the MT2ST model notably outperforms conventional embedding training methods in terms of training efficiency and overall quality. The newly developed embeddings exhibit superior performance and robustness across a range of downstream NLP tasks, including sentiment analysis, part-of-speech tagging, and named entity recognition. These results underscore the effectiveness of MT2ST in reducing training times and lessening computational resource requirements. The multi-task to single task learning also has applications in some recent work, such as [WCL24] [LWC+24] [XQW+24] [DLL24]. This efficiency makes MT2ST a valuable solution for fast training and high-quality word representations.

### v3

@@ Abstract
Efficient machine learning (ML) has become increasingly important as models grow larger and data volumes expand. In this work, we address the trade-off between generalization in multi-task learning (MTL) and precision in single-task learning (STL) by introducing the Multi-Task to Single-Task (MT2ST) framework. MT2ST is designed to enhance training efficiency and accuracy in word embedding tasks, showcasing its value as a practical application of efficient ML.
Our framework employs two strategies: *Diminish*, which gradually reduces the influence of auxiliary tasks, and *Switch*, which transitions training from MTL to STL at a specific point. Empirical results show that MT2ST reduces training time by 67% compared to STL and by 13% compared to traditional MTL, while maintaining high accuracy. These findings highlight MT2ST as an efficient ML solution tailored for optimizing word embedding training. Code is available at https://github.com/NoakLiu/MT2ST.
@@
Efficient ML, Model Compression, ML Acceleration
@@ I Introduction
Efficient machine learning (ML) has become increasingly crucial due to the growing size of models and datasets across various domains such as natural language processing (NLP), computer vision, and speech recognition. The need for efficient ML has driven innovations in areas such as embedding generation for semantic representation, multi-task learning (MTL) for generalization across related tasks, and optimization techniques that reduce computational demands. However, the training of embeddings faces several challenges, including computational inefficiencies, optimization difficulties, and scalability issues, as highlighted by Mikolov et al. (2013) [MCCD13] and more recent surveys on NLP efficiency [Z+23].
Embedding training typically utilizes either single-task learning (STL) or multi-task learning (MTL). STL is known for providing precise task-specific learning but lacks the broader generalization capabilities of MTL, as discussed by Collobert et al. (2011) [CW11]. On the other hand, MTL enables models to share representations across tasks, enhancing generalization. However, it introduces challenges such as gradient conflicts and inefficiencies in resource usage, particularly when dealing with tasks that require distinct representations [SK18, Z+23]. This makes balancing the breadth of MTL with the precision of STL an ongoing challenge in machine learning.
In this paper, we introduce the Multi-Task to Single-Task (MT2ST) framework, which addresses this challenge by combining the strengths of both MTL and STL. MT2ST employs two novel strategies: Diminish and Switch. The Diminish strategy gradually reduces the influence of auxiliary tasks over time, ensuring a smooth transition to STL, while the Switch strategy shifts the model from multi-tasking to single-tasking at a predefined point. Our framework significantly improves training efficiency and accuracy, as demonstrated in our empirical studies. Specifically, we show that MT2ST reduces training time by 67% compared to STL and by 13% compared to traditional MTL, while maintaining high accuracy and model performance.
The contributions of our work can be summarized as follows:
  • Bridging the gap between MTL and STL with the innovative MT2ST framework.
  • Introducing the Diminish and Switch strategies for efficient learning.
  • Demonstrating MT2ST’s superior training efficiency and performance in word embedding tasks.
@@ II Motivation
II-A Challenges in Single-Task Embedding Training
Embeddings are widely used for representing objects, such as words, images, or even users, in continuous vector spaces that capture their semantic relationships. Despite their broad applicability across domains, embedding generation faces several challenges:
@@ II-A Challenges in Single-Task Embedding Training
Data Volume: Large-scale datasets are often required for training high-quality embeddings, resulting in significant time and computational resource consumption. Processing these large datasets is inherently slow and computationally expensive [EWVD19, LJP24].
Computational Complexity: The generation of embeddings involves complex mathematical computations, demanding high computational power and time. Efficient computational methods are essential to optimize these processes. Recent work has focused on algorithmic optimizations and system co-design to reduce computational demands, such as in [LWJV24, LJP24].
Optimization Difficulties: Training embeddings for accurate semantic or contextual representation is a complex process, requiring iterative optimization. This optimization is resource-intensive and time-consuming, particularly when dealing with large datasets or high-dimensional embeddings. Research such as [G+21, S+21] highlights the challenges in optimizing embeddings while maintaining their quality across different tasks.
II-B Optimizing Training Efficiency with Multi-Task Learning
Embedding generation can be made more efficient by using multi-task learning (MTL), where models are trained on related tasks simultaneously. This allows for the sharing of representations across tasks, enhancing the quality of the embeddings by leveraging additional, related information. MTL has been widely applied in various domains, such as NLP, computer vision, and speech recognition. For example, in computer vision, multi-task networks improve object detection and segmentation by learning shared features across related tasks [M+16, Z+20]. In NLP, MTL improves contextual understanding of embeddings [ZYY+23, SSM+22].
Recent work has shown that MTL can significantly improve efficiency, not only in NLP but also in other domains like computer vision or speech recognition, where it is applied to tasks such as object detection or speaker identification. MTL enables models to generalize across multiple tasks, leading to more robust representations [G+21].
II-C Inefficiencies in Multi-Task Learning for Embedding Generation
Despite its benefits, MTL can also introduce inefficiencies, especially in the context of embedding generation. One issue arises when different tasks impose conflicting gradients on the shared embedding representation. For instance, while one task might prioritize fine-grained details, another might emphasize broader patterns, leading to conflicting gradient updates that can hinder the optimization process. This conflict is further discussed in recent works such as [LLJ+21].
Furthermore, MTL can be computationally expensive, as each task requires updates to the shared embedding layer, resulting in increased resource consumption and longer training times. This is particularly problematic when working with large-scale datasets or high-dimensional embeddings. Work by [liu2020multi] and [SSM+22] has examined the computational costs and the challenges that arise with increasing task complexity.
@@ III-A MT2ST Framework
We introduce the MT2ST (Multi-Task to Single-Task) framework to optimize embedding generation training. It combines multi-task learning (MTL) and single-task learning (STL) to achieve efficient training while overcoming common challenges in multi-task environments.
The process starts with MTL, where a unified model with a shared embedding layer is trained across multiple tasks. This allows the model to capture diverse linguistic features and semantic knowledge. The shared embedding layer benefits from varied inputs, providing a more generalized word representation [LJD19].
After the MTL phase, MT2ST transitions to STL, fine-tuning the pre-trained embeddings for specific tasks. This phase refines the embeddings to match the unique requirements of each task, improving performance while retaining the knowledge gained from the MTL phase. Techniques like adaptive learning rates and selective freezing of embedding dimensions ensure a smooth transition and maintain the balance between generalization and specialization [TLJ+23].
@@ III-B Model Construction
Instead of mono-task fine-tuning used in Multi-Task Supervised Pretraining [MST+20], our MT2ST model develop an adaptive method to accelerate embedding generation through a multi-task to single-task learning framework.
@@ III-B2 Embedding Layer
  • Learns dense vector representations of embedding, capturing semantic and syntactic information.
@@ III-C Model Overview
Figure 1: MT2ST Embedding Training Framework Overview
@@ III-E MT2ST: Switch Strategy
Post $T_{switch}$, the update focuses solely on the primary task:
@@ IV Result Analysis
We conducted several experiments to evaluate the performance of the MT2ST framework, focusing on model compression, training time, and convergence speed. The experimental details are provided in the appendix.
IV-A Baseline (Single-task Fine-Tuning)
In the first experimental setup, we fine-tuned pre-trained word embeddings on a single task dataset. While the model showed a steady decrease in loss, we observed fluctuating accuracy, indicating overfitting and slow convergence. Despite achieving relatively high accuracy, the model’s training time and memory usage were high due to the lack of task-sharing strategies.
Model Setup Acc (%) Training Time (s) Compression Rate (%)
ST Fine-tuning 68 108 0%
TABLE I: Results of the Single-task Fine-Tuning Experiment
IV-B Multi-task Learning (MTL)
In the second experimental series, we applied a multi-task learning (MTL) approach. This setup allowed the model to share the embedding layer across tasks, improving efficiency by leveraging a broad set of linguistic features. As expected, the model exhibited a smoother loss decrease and a more stable convergence.
Model Setup Acc (%) Training Time (s) Compression Rate (%)
MTL 68 60 21%
TABLE II: Results of the Multi-task Learning Experiment
However, MTL still faced challenges in training time due to conflicting gradients from tasks with divergent goals. The training time was slightly reduced, but the compression rate only improved by 21%.
IV-C MT2ST: Diminish Strategy
In the third experiment, we employed the MT2ST-diminish strategy. This strategy gradually reduces the influence of auxiliary tasks over time, ensuring a smooth transition to single-task learning. We observed a more efficient convergence with a final test accuracy of 71%, surpassing the MTL approach. The model achieved a significantly better compression rate and reduced training time, achieving faster convergence and maintaining strong performance on the test set.
Model Setup Acc (%) Training Time (s) Compression Rate (%)
MT2ST: Diminish 71 72 44%
TABLE III: Results of the MT2ST-Diminish Strategy Experiment
Compared to the MTL approach, the MT2ST-diminish strategy reduced training time by 20% and increased the compression rate by 44%, showing that it effectively balances efficiency and performance.
IV-D MT2ST: Switch Strategy
Finally, we implemented the MT2ST-switch strategy, which introduces a more abrupt transition between multi-task and single-task learning phases. The model achieved a final accuracy of 77% on the test set, with a dramatic improvement in convergence speed. Notably, the loss function showed a sharp drop after the switch, indicating efficient transition and adaptation to the task-specific learning phase.
Model Setup Acc (%) Training Time (s) Compression Rate (%)
MT2ST: Switch 77 64.8 53%
TABLE IV: Results of the MT2ST-Switch Strategy Experiment
The MT2ST-switch strategy achieved a 53% compression rate, reducing the training time by 40% compared to MTL while also improving accuracy. This demonstrates that MT2ST not only accelerates convergence but also optimizes computational efficiency.
IV-E Summary of Results
The experimental results show that MT2ST strategies, particularly the Diminish and Switch strategies, offer significant improvements in both model compression and training time reduction. The following table summarizes the overall results:
Model Setup Acc (%) Training Time (s) Compression Rate (%)
Single-task Fine-tuning 68 108 0%
Multi-task Learning (MTL) 68 60 21%
MT2ST: Diminish 71 72 44%
MT2ST: Switch 77 64.8 53%
TABLE V: Summary of All Experimental Results
These results confirm the effectiveness of the MT2ST framework in improving both computational efficiency and convergence speed while maintaining high accuracy across tasks.
@@ V Discussion
The experimental results highlight the effectiveness of the MT2ST framework in improving both model compression and accelerating training process. Specifically, the MT2ST-Switch strategy achieves a 53% compression rate and reduces training time by 40%, while still maintaining 77% accuracy. This outcome confirms that the MT2ST framework offers a significant computational advantage, demonstrating that moderate losses in accuracy are acceptable when balanced against substantial resource savings. Such trade-offs are consistent with previous work on model compression, where slight performance degradation occurs to achieve higher efficiency [HPTD15].
In comparison to the baseline single-task fine-tuning approach, multi-task learning (MTL) leads to a compression rate of 21% and a 45% reduction in training time, but still suffers from slower convergence and increased computational demands due to conflicting gradients across tasks. This suggests that while MTL benefits from sharing task-specific knowledge, it can face challenges in balancing efficiency and accuracy across diverse tasks [Rud17]. The MT2ST-Diminish strategy, which gradually reduces the influence of auxiliary tasks, provides a more efficient approach, achieving a 44% compression rate while maintaining a relatively high accuracy of 71%. This result supports the effectiveness of gradual task reduction in enhancing both performance and resource utilization [ZCZ20].
Another key insight from these experiments is the balance between accuracy and model compression. As expected, compression strategies such as pruning and quantization lead to minor drops in accuracy. However, the gains in computational efficiency often outweigh these losses, especially in resource-constrained environments like edge devices or mobile applications [LHW20]. The MT2ST framework, by combining multi-task and single-task learning in a smooth transition, appears well-suited for real-world applications where training time and memory usage are critical.
Furthermore, the decrease in accuracy observed with higher compression rates is a common phenomenon in deep learning model optimization, as demonstrated by previous studies on network pruning [MAV17]. This phenomenon indicates that, while compression techniques reduce model size, they must be carefully calibrated to avoid overfitting or underfitting. Overall, the MT2ST framework provides a promising solution for the computational resources reduction and training acceleration.
Appendix A Experimental Results Figures
This section includes the figures corresponding to the experimental results presented in the main text.
A-A Single-task Fine-Tuning
Figure 2 shows the loss and accuracy changes for the single-task fine-tuning approach.
Figure 2: Loss and Accuracy Change for Single-task Fine-Tuning
A-B Multi-task Learning (MTL)
Figures 3 and 4 show the loss and accuracy changes for the multi-task learning approach.
Figure 3: Loss Change in Multi-task Learning
Figure 4: Accuracy Change in Multi-task Learning
A-C MT2ST: Diminish Strategy
Figures 5 and 6 show the loss and accuracy changes for the MT2ST-diminish strategy.
Figure 5: Loss Change in MT2ST: Diminish Strategy
Figure 6: Accuracy Change in MT2ST: Diminish Strategy
A-D MT2ST: Switch Strategy
Figures 7 and 8 show the loss and accuracy changes for the MT2ST-switch strategy.
Figure 7: Loss Change in MT2ST: Switch Strategy
Figure 8: Accuracy Change in MT2ST: Switch Strategy